OpenStudy (anonymous):

What is the equation of the line that goes through (-1,-1) and is perpendicular to y=5x+3

OpenStudy (anonymous):

y= mx+3 can u tell me the slope of this line?

OpenStudy (anonymous):

Wouldn't the slope be -1/5? :3

OpenStudy (anonymous):

u need to find the equation of a line, which is perpendicular to y=5x+3 what is the slope of the given line? can u draw this line on x-y plane?

OpenStudy (anonymous):

1st draw this line...y=5x+3

OpenStudy (anonymous):

I never got taught how to do it like that :( But I think the answer is y+1=-1/5(x+1) Correct me if I'm wrong please!

OpenStudy (anonymous):

do u know that y=5x+3 represents a line?

OpenStudy (anonymous):

u want answer or learn to solve this?

OpenStudy (anonymous):

I have a feeling I know how to solve it. I just can't think straight right now lol.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is my answer right tho?

OpenStudy (anonymous):

but u want answer...right?

OpenStudy (anonymous):

Yes, please

OpenStudy (anonymous):

yes...ur answer is right..:)

OpenStudy (anonymous):

Yay, ok thanks !! :)

OpenStudy (anonymous):

:)